Write an equivalent fraction for:
a. 1/4.
b. 1/2.
c. 6/8.
d. 2/4.
e. 2/2.
f. 2/8.
g. 8/8.
h. 4/8.
i. 4/4.
1
Expert's answer
2020-09-23T17:29:06-0400

a. 1/4.

b. 1/2.

c. 6/8=3/4.

d. 2/4=1/2.

e. 2/2=1.

f. 2/8=1/4.

g. 8/8=1.

h. 4/8=1/2.

i. 4/4=1.


a=f

b=d=h

e=g=i